OpenStudy (anonymous):

When you have a power outside parentheses and a power inside, does it multiply them together or add them? EX: (3x^-3/4x^3)^-2

OpenStudy (amistre64):

hmm\[(a^3)^2=a^3*a^3=a^{(3+3)}\]

OpenStudy (amistre64):

so, if 3+3 = 5, it adds them if 3+3 = 6, it multiplies them